# How Linguistics Learned to Stop Worrying and Love the Language Models

### v1

@@ B. Matching verb, varying context
Of course, this work has not all been positive in its results (for example, Weissweiler et al., 2023b, ), and it is possible for models to be “right for the wrong reasons” (McCoy et al.,, 2019), adopting shallow heuristics which make correct predictions on certain test sets but do not generalize properly, or models may show mastery of linguistic form without a concomitant ability to understand the implications of utterances (Weissweiler et al.,, 2022)—in effect mastering linguistic form without mastering linguistic function (Mahowald et al.,, 2024). As an example of how shallow heuristics may lead to the illusion of good performance, Vázquez Martínez et al., (2023) find that a simple $n$-gram model performs surprisingly well on the paired grammatical and ungrammatical sentences of the BLiMP dataset (Warstadt et al.,, 2020), which turn out to be poorly controlled for $n$-gram frequency. [^4] We note that this criticism does not apply to studies based on more carefully selected sets of sentences (for example, Marvin and Linzen,, 2018; Futrell et al., 2019b, ; Hu et al.,, 2020; Wilcox et al., 2023a, ).
@@
The possibility that models generalize on the basis of shallow heuristics motivates deeper investigation. We mention two approaches here. The first involves controlling the training data for language models and then observing their generalizations on evaluation data that is unlike anything in their training data. This approach allows one to study true generalization in language models, rather than shallow memorized patterns. For instance, Misra and Mahowald, (2024) found that the relatively rare A+Adjective+Numeral+Noun construction “a beautiful five days in Texas” is learned to some extent by models in which instances of the construction are systematically withheld during training—suggesting generalization. Ahuja et al., (2024), building on McCoy et al., (2020), train Transformer language models on a corpus of English-like text that has been constrained so that subjects and verbs are always adjacent. That is, the corpus contains sentences like I saw the key to the cabinets, but never something like The key to the cabinets is on the table. The question is, then, if the model trained in this way will generalize in the way that humans do, preferring The key to the cabinets is on the table (where the form of the verb depends on the subject of the sentence) over something like The key to the cabinets are on the table (where the form of the verb depends on the linearly previous verb, which is perfectly consistent with the training data). Ahuja et al., (2024) find that Transformer language models do make the human-like generalization, showing that they can learn the formal structure underlying nonlocal subject–verb agreement without ever observing nonlocal subject–verb agreement (see also Jumelet and Zuidema,, 2023; Patil et al.,, 2024).
@@
Another approach is to dig into the language models’ internals, using approaches such as probing, where one attempts to decode linguistic representations from the internal representations of neural networks, and causal interventions, where model internals are changed and the resulting output changes are observed. Such studies have increasingly revealed genuine representations of grammatical structure once neural networks’ internal representations are decoded properly (Lakretz et al.,, 2019; Voita and Titov,, 2020; Manning et al.,, 2020; Papadimitriou et al.,, 2021; Ravfogel et al.,, 2021; Lampinen,, 2024; Diego-Simón et al.,, 2024, among others). For example, Hewitt and Manning, (2019) and Chi et al., (2020) find that syntactic relationships among words in sentences are encoded geometrically in a linear subspace of the high-dimensional contextual embedding space for words. Ravfogel et al., (2021) showed that these syntactic representations are causally active: systematically altering them changes model behavior in the ways we would expect.
@@
An important caveat to these findings is that the various evaluations of structural representation in language models have mostly used English as the target language (Blasi et al.,, 2022). As such, current architectures and paradigms are likely biased towards English and other widely spoken languages. An important area for future work is to extend the empirical scope of these studies to other languages whose formal patterns may differ from English (Bender,, 2022).
@@
There are ongoing debates about the extent to which language models capture more complex formal patterns within languages (Vázquez Martínez et al.,, 2023; Lan et al.,, 2024) or to what extent they can be said to “understand” (Bender and Koller,, 2020) or refer to things in the world (Mandelkern and Linzen,, 2023; Lederman and Mahowald,, 2024). This is not the right place to adjudicate these debates. What we think is uncontroversial, based on the work summarized above, is that language models have learned nontrivial formal patterns corresponding to linguistic structure, to a greater extent than many expected was possible. Which is to say, we believe that even if language models are not perfect at grammar, they nevertheless really have learned something about linguistic structure, in a way that problematizes earlier claims about the role of statistics in language learning. They are not purely relying on heuristics or memorization or cheap tricks. They have, to a meaningful extent, learned “the real thing”—that is, the thing that we care about, as language scientists interested in questions like how languages are learned, how they are processed, how and why they vary, and where they come from.
@@
We now proceed to ask why this is interesting for the science of language, and how the science of language productively integrate these insights.
@@ 3 The success of LMs is interesting for the science of language
“…nor did Pnin, as a teacher, ever presume to approach the loft halls of modern scientific linguistics, that ascetic fraternity of phonemes, that temple wherein young people are taught not the language itself, but the method of teaching others to teach that method; which method, like a waterfall splashing from rock to rock, ceases to be a medium of rational navigation but perhaps in some fabulous future may become instrumental in evolving esoteric dialects—Basic Basque and so forth—spoken only by certain elaborate machines.”
@@
LMs learn nontrivial aspects of language. But one could still object that this result is not relevant to language science because human language is the product of the human brain (and/or vocal tract, body, social environment, etc.), whereas LMs have a totally different architecture and different objective. Perhaps these differences are so great that, just as there is limited value to the ichthyologist studying a submarine, a language scientist has nothing to gain from studying language models.
Here we argue that the success of language models is relevant, indeed important, to the science of language, because (1) there are good reasons to believe that there should be parallels between engineering artifacts like language models and human language, and (2) the success of language models in learning from text upends ways of thinking that are deeply ingrained in generative linguistics and parts of cognitive science, and (3) language models are already connected to statistical and probabilistic traditions in linguistics, and in fact arose out of them to a large extent. While some ascribe the success of LMs to engineering alone, LMs’ success is also a success of an intellectual tradition within linguistics. As such, LMs aren’t alien invaders into linguistics from engineering. Rather, they are tools of the kind that have long been used to answer fundamentally linguistic questions.
@@ 3.1 Parallels between engineering models and cognition
From the study of vision, there is strong precedent for the idea that neural networks developed purely for practical applications can tell us a great deal about cognition as it is implemented in the brain. Similarly, we think there is good reason to think that LMs will be informative for the study of human language.
@@
Hubel and Wiesel, (1959) discovered that early processing of visual information is performed by neurons that are selectively responsive to edges. After this discovery, the question became why early visual processing works this way. Olshausen and Field, (1996) showed that edge filters resembling those found in the visual cortex were the generic solution to the problem of representing visual information accurately in a neural network under a constraint that only a small number of units should be active (that is, a sparsity constraint, arising ultimately from a power constraint on neural firing). Even better explanation came from the engineering of artificial visual systems, in particular the development of AlexNet, a large (for the time) hierarchical convolutional neural network for image classification (Krizhevsky et al.,, 2012). Yamins et al., (2014) showed that this architecture, when trained to do object recognition using the same engineering methods and techniques as in the machine learning model, not only developed edge filters in its early layers, but also receptive fields in later layers corresponding to later layers of visual processing in the primate brain. The overall picture that emerges from this line of work is: the neuronal organization of visual processing is determined by the function of finding a sparse code to identify and manipulate objects in the environment, a function largely shared between the biological and artificial systems.
@@
Here, the artificial system provided guiding insights for understanding the natural system. Something similar is happening in the science of language processing, where the internal representations developed by language models are predictive of activation patterns in the human language network (Goldstein et al.,, 2022; Rathi et al.,, 2024). The picture is not yet as clear as it is in vision—this is an area of active research, and there are no animal models we can use to understand language in the brain—but there is precedent to think that the artificial system will be informative about the natural system here too.
@@
Why did this work in vision? To explain the success of the approach, Cao and Yamins, (2021) introduce the Contravariance Principle: in short, if a computational problem is hard, then there are only a few ways to solve it, so we expect different systems (for example, neural networks and the brain) to converge to the same solution (see also Huh et al.,, 2024). Vision is hard, therefore we can expect that any system that solves the problem will share core features with any other. We see this not only in the comparison of humans vs. neural networks, but also the comparison of humans vs. monkeys (Rajalingham et al.,, 2018), and primates vs. animals whose visual cortices arise from totally different evolutionary phylogenies, such as cephalopods (Pungor et al.,, 2023).
@@
It seems undeniable that learning and using language is hard, and as we have discussed, large Transformer-based models do learn and deploy nontrivial aspects of language. If the Contravariance Principle is right, then it is reasonable to expect some degree of commonality in the solutions that exist in the brain and in Transformers. Indeed, the massive literature on linguistic interpretability in Transformers, which we partially reviewed in Section 2.2, has revealed representational strategies in neural networks that lead to correct novel predictions about human performance (Lakretz et al.,, 2021), and the way that Transformers process syntactic features like agreement cues has intriguing parallels with independently-proposed cognitive frameworks for modeling human language processing (Lewis and Vasishth,, 2005; Ryu and Lewis,, 2021; Timkey and Linzen,, 2023).
@@
It is just not the case that these models can be dismissed out of hand because they are different from the brain. To do so would be to miss out on a source of in silico insight and inspiration which has been hugely useful in other areas of cognition.
@@ 3.2.1 The significance of the learning problem in linguistics and cognitive science
For a century, cognitive scientists, linguists, computer scientists, and philosophers developed formal theories of how language learning—or any learning—must work. The key concept, known under many names, has been inductive bias (Goyal and Bengio,, 2022), as illustrated in a few ways in Figure 3.2.1. We can think of a learner as a device that takes in some data and outputs a hypothesis, or a set of hypotheses, or a probability distribution over hypotheses, for the underlying process generating the data. For example, given a bunch of data points in a 2D plane, one hypothesis would be that the points are generated along a line in the plane and then perturbed by noise; this is the assumption underlying linear regression. Given linguistic utterances as data, one might hypothesize a certain grammar underlying the data.
The key issue is that, in a very general sense, the data fundamentally underdetermine the hypothesis that the learner arrives at, because any data is compatible with many, many hypotheses. In Figure 3.2.1A, where the learning data is points and the hypotheses are lines, both the straight and the curvy lines fit the learning data equally well; the fact that the straight line seems like the right hypothesis—and is in fact the one that humans would seize on—therefore cannot be a function of fit to the data. It must be a function of fit to the data plus something that biases a learner to favor one hypothesis over another, among hypotheses that fit the data equally well: this is inductive bias. Inductive bias is something that the model (or modeler) brings to the problem, not something inherent in the data.
@@ B. Linguistic data
(a) Illustration of the role of inductive bias in generalization. A. The black points represent data, and the lines represent generalizations that could be formed on the basis of the data. Both lines fit the data exactly, but make different predictions for new datapoints. Intuitively, the blue line seems like a better hypothesis. But the choice to prefer any single generalization over any other can only be a function of inductive bias—a preference for hypotheses which are in some sense ‘simpler’—which is not a function of the data. B. Inductive bias in language. In black, possible corpus data from English. In blue, sentences which generalize from the corpus data on the basis of hierarchical structure. In red, sentences which generalize on the basis of linear word order. Intuitively, the blue sentences seem like natural generalizations and the red ones seem unnatural, even though all are equally consistent with the data given. This is claimed to be the result of innate inductive biases (Chomsky,, 1971).
@@
Inductive bias shows up in a number of forms and under many names in the scientific literature. In linguistic theory, it appears perhaps first as the evaluation procedure in Chomsky, (1965, pp. 30–48), a function which compares two different grammars that can generate an observed set of sentences (that is, two grammars that fit some training data equally well), and ranks them in order of preference, in order to model how humans generalize beyond the data (Chomsky,, 1965, p. 45). More generally, it appears as Universal Grammar—an “innate schematism of mind that is applied to the data of experience” (Chomsky,, 1971, p. 28) to enable language learning and generalization. In this approach, (generative) grammatical theories can instantiate hypotheses about that schematism; they delimit what languages may exist and be learned from data.
@@
This approach to linguistic theory serves the goal of explanatory adequacy: the idea is that a theory of grammar should not only capture which sentences are grammatical and ungrammatical in a particular language, but also that the theory itself should predict the generalizations that people make about language, and the languages that we actually find in the world (Chomsky,, 1965, Ch. 1). Thus if a grammatical theory seems to accommodate ‘impossible’ or unattested kinds of languages, then this would be considered evidence against that theory, because the theory is supposed to represent the “schematism of mind” or inductive bias that a learner uses to generalize from input data. This approach to explaining linguistic phenomena has been called “explanation by constrained description” (Haspelmath,, 2009, pp. 384–385).
@@
Naturally then, inductive bias has been the major focus of study in language learning, seen as the simultaneous solution to two different problems: (1) how children can learn language from inadequate data (the classic Argument from the Poverty of the Stimulus: see Pearl,, 2022), and (2) why human language is the way it is: because if a language went against the child’s inductive biases, it would be hard or impossible for the child to learn. [^5] Given the centrality of this reasoning to much of linguistic theory, one might expect it to be backed up with strong experimental evidence that humans cannot or do not learn languages which violate the putatively universal principles of human languages. This is not the case: there is only limited and ambiguous experimental evidence for hard formal limits on human language learning. Among the most on-target existing studies is Smith et al., (1993), whose object of study was a man described as a polyglot savant living in a mental health facility. This individual and four control subjects (linguistics undergraduates) were tasked with learning artificial languages designed to be ‘impossible’ in three ways: (1) negation and tense are indicated by word order, (2) there is an agreement pattern judged to be impossible, and (3) the position of an emphatic marker is determined by a rule involving counting words. Results are not systematically reported, but seem to indicate that the polyglot was able to learn the ‘impossible’ word order and agreement rules (1) and (2), but not the rule for the emphatic marker (3). Another commonly cited experimental work on this topic is Musso et al., (2003), who expose German speakers to Italian and Japanese sentences, either following the real rules of those languages, or following modified rules deemed to be linguistically impossible, for example placing a negation marker after the third morpheme from the beginning of a sentence. The result is equally accurate learning of the ‘natural’ and ‘unnatural’ languages. fMRI on the subjects shows that the real languages elicit activity in the left inferior frontal gyrus, while the unnatural ones elicit activity elsewhere. We believe the meaning of these results is unclear. Only a small number of languages and participants (all of whom were already native speakers of largely hierarchically-structured languages) were tested, language localization in the brain is still contentious, and the patterns of brain activity for the ‘unnatural’ languages might reflect a lack of practice with such patterns, rather than their impossibility or a qualitative difference between linear and hierarchical rules. The dream has been to come up with a formalism for linguistic description which also captures human generalizations—a Universal Grammar, corresponding to a genetic endowment unique to humans and domain-specific to language, and perhaps reflecting representational structures in the brain (Chomsky,, 1988; Hauser et al.,, 2002; Berwick et al.,, 2011)—thus in one fell swoop solving (1) and (2) and also explaining why only humans have language.
@@
Under this logic, it follows that a good theory of language learning should be restrictive: that is, there should be languages that cannot be learned under the theory, and this restriction on the hypothesis space provides explanatory adequacy. [^6] See, for example, Kodner et al.,’s (2022) criticism of Yang and Piantadosi,’s (2022) model of language learning as Bayesian program induction, a model which successfully learns grammars of various formal classes given small amounts of string input, thus addressing the Poverty of the Stimulus problem. Although the model readily meets the challenge of inducing formal structure from strings, it has been dismissed by some in the linguistics literature because the same model could also learn grammars that are unlike human language. In terms of Figure 3.2.1, the intuitively bad hypotheses would be ruled out as simply unavailable as mental representations during learning. This approach to linguistic explanation has a pleasing elegance to it: learners must be restricted to learn properly, and we see that the variation in actual languages is restricted, therefore we can kill two birds with one stone by finding the right representational restrictions on learners. [^7] For a modern and formally fleshed out take on this approach, see the Subregular Program in phonology (Heinz,, 2018) (and syntax and semantics). A group of formal linguists have argued that phonological generalizations fall into various formal classes called subregular. These formal classes have strong formal guarantees on learning from finite data when the learner is assumed to be restricted to a subregular hypothesis space.

### v2

@@ B. Matching verb, varying context
Of course, this work has not all been positive in its results. Models can be “right for the wrong reasons” (McCoy et al.,, 2019), adopting shallow heuristics which make correct predictions on certain test sets but do not generalize properly, or models may show mastery of linguistic form without a concomitant ability to understand the implications of utterances (Weissweiler et al.,, 2022)—in effect mastering linguistic form without mastering linguistic function (Mahowald et al.,, 2024). As an example of how shallow heuristics may lead to the illusion of good performance, a simple $n$-gram model performs fairly well on some subsets of paired grammatical and ungrammatical sentences of the BLiMP dataset (Warstadt et al.,, 2020) (although not as well as a neural LM), which does not explicitly control for $n$-gram frequency (Vázquez Martínez et al.,, 2023). [^4] We note that these kinds of criticisms are harder to apply to studies based on more carefully controlled sets of sentences (for example, Marvin and Linzen,, 2018; Futrell et al., 2019b, ; Hu et al.,, 2020; Wilcox et al., 2023a, ).
@@
The possibility that models generalize on the basis of shallow heuristics motivates deeper investigation. We discuss two approaches here. The first involves controlling a model’s training data and then observing its generalizations on evaluation data that is unlike anything in the training data (Jumelet et al.,, 2021; Feng et al., 2024b, ; Misra and Mahowald,, 2024; Leong and Linzen,, 2023; Yao et al.,, 2025). This approach allows one to study true generalization in language models, rather than shallow memorized patterns. Ahuja et al., (2025) trained Transformer language models on a corpus of English-like text that has been constrained so that subjects and verbs are always adjacent. That is, the corpus contains sentences like I saw the key to the cabinets, but never something like The key to the cabinets is on the table. The question is, then, if the model trained in this way will generalize in the way that humans do, preferring The key to the cabinets is on the table (where the form of the verb depends on the subject of the sentence) over something like The key to the cabinets are on the table (where the form of the verb depends on the linearly previous verb, which is perfectly consistent with the training data). Ahuja et al., (2025) find that neural language models do make the human-like generalization, providing evidence that they can learn the formal structure underlying nonlocal subject–verb agreement without ever observing nonlocal subject–verb agreement (see also Patil et al.,, 2024).
@@
Another approach is to dig into the language models’ internal states. Large language models have a reputation of being black boxes, whose internal processes and representations are inscrutable, but researchers have been gaining growing traction on understanding their inner workings. Approaches include probing, where one attempts to decode linguistic features from the internal representations of neural networks, and causal interventions, where model internals are changed and the resulting output changes are observed (Hewitt and Manning,, 2019; Chi et al.,, 2020; Voita and Titov,, 2020; Manning et al.,, 2020; Papadimitriou et al.,, 2021; Ravfogel et al.,, 2021; Lampinen,, 2024; Diego-Simón et al.,, 2024, among others). Such studies have revealed increasing evidence of grammatical structure in models. We consider these kinds of studies, particularly ones that probe models causally, to be promising avenues for linguistics and cognitive science (see Section 4.2).
@@
A caveat to these findings is that the various evaluations of structural representation in language models have mostly used English (or a handful of other languages) as the target (Blasi et al.,, 2022). An important area for future work is to extend the empirical scope of these studies to a wider set of languages (work which is ongoing; see Jumelet et al.,, 2025, for a multilingual grammatical benchmark).
@@
While there is disagreement about how much language models capture more complex formal patterns (Vázquez Martínez et al.,, 2023; Lan et al.,, 2024; Someya et al.,, 2024) or to what extent they can be said to “understand” (Bender and Koller,, 2020) or refer to things in the world (Mandelkern and Linzen,, 2024; Lederman and Mahowald,, 2024), this is not the right place to adjudicate these debates. What is clear is that language models have learned nontrivial formal linguistic patterns better than many expected was possible. That is to say, they really have learned something about linguistic structure, in a way that problematizes earlier claims about the role of statistics in language learning. They are not purely relying on heuristics or memorization or cheap tricks. They have, to a meaningful extent, learned “the real thing”—that is, the thing that we care about, as language scientists interested in questions like how languages are learned, how they are processed, how and why they vary, and where they come from.
@@
We next ask why this is interesting for the science of language, and how the science of language productively integrates these insights.
@@ 3 The success of LMs is interesting for the science of language
“…nor did Pnin, as a teacher, ever presume to approach the lofty halls of modern scientific linguistics, that ascetic fraternity of phonemes, that temple wherein young people are taught not the language itself, but the method of teaching others to teach that method; which method, like a waterfall splashing from rock to rock, ceases to be a medium of rational navigation but perhaps in some fabulous future may become instrumental in evolving esoteric dialects—Basic Basque and so forth—spoken only by certain elaborate machines.”
@@
LMs learn nontrivial aspects of language. But one could still object that this result is not relevant to language science because human language is the product of the human brain (and/or vocal tract, body, social environment, or what have you), whereas LMs have a totally different architecture and different objective. Perhaps these differences are so great that a language scientist has practically nothing to gain from studying language models, in the same way that there is limited value to an ornithologist studying jet engines, even though both bird wings and jet engines enable flight (Kodner et al.,, 2023).
Here we argue that the success of language models is relevant, indeed important, to the science of language, because (1) there are good reasons to believe that there should be parallels between engineering artifacts like language models and human language, and (2) the success of language models in learning from text upends ways of thinking that are deeply ingrained in generative linguistics and parts of cognitive science, and (3) language models are already connected to statistical and probabilistic traditions in linguistics, and in fact arose out of them to a large extent. As such, LMs aren’t alien invaders into linguistics from engineering. Rather, they are tools similar to those that have long been used to answer fundamentally linguistic questions.
@@ 3.1 Parallels between engineering models and cognition
From the study of vision, there is strong precedent for the idea that neural networks developed purely for practical applications can tell us a great deal about cognition as it is implemented in the brain. Similarly, there is good reason to think that LMs will be informative for the study of human language. To elucidate this example, we briefly review, in simplified form, the development of the neuroscience of vision and the role of neural networks in it.
@@
Hubel and Wiesel, (1959) discovered that early processing of visual information is performed by neurons that are selectively responsive to edges in the visual input. After this discovery, the question became why early visual processing works this way. Olshausen and Field, (1996), building on information-theory-inspired intuitions from Barlow, (1961, 1989), showed that edge detectors resembling those found in the visual cortex were the generic solution to the problem of representing visual information accurately in a neural network under a constraint that only a small number of units should be active (that is, a sparsity constraint, arising ultimately from a power constraint on neural firing). Even better explanation—both in the sense of explanatory depth and predictive accuracy—came from the engineering of artificial visual systems, in particular the development of AlexNet, a large (for the time) hierarchical convolutional neural network for image classification (Krizhevsky et al.,, 2012). Yamins et al., (2014) showed that this architecture, when trained to do object recognition, not only developed edge filters in its early layers, but also receptive fields in later layers corresponding to later layers of visual processing in the primate brain. The overall picture that emerges from this line of work is: the neuronal organization of visual processing is determined by the function of finding a sparse code to identify and manipulate objects in the environment, a function largely shared between biological and artificial systems.
@@
Here, the artificial system provided guiding insights for understanding the natural system. [^5] There remains controversy over the extent to which deep neural networks are the best model of human vision (see Bowers et al.,, 2023, for a more skeptical take along with spirited replies), but even skeptics of neural models judge that this research program has been fruitful. Something similar is now happening in the science of language processing, where the internal representations developed by language models are predictive of activation patterns in language areas of the brain (Goldstein et al.,, 2022; Caucheteux et al.,, 2023; Hosseini et al., 2024b, ; Rathi et al.,, 2024), and predictability as estimated by a language model is an important factor in studies that predict neural activity (Stanojević et al.,, 2023; Zhao et al.,, 2025). The picture is not yet as clear as it is in vision—this is an area of active research, and there are no animal models we can use to get the plentiful high-resolution controlled neural data we would like—but there is precedent to think that the artificial system will be informative about the natural system here too.
Why did this work in vision? To explain the success of the approach, Cao and Yamins, (2021) introduce the Contravariance Principle: the idea that, if we want to uncover solutions to problems that are common between brains and models, we should focus on hard problems. The reason is that, if a computational problem is hard in the sense that it requires satsifying multiple potentially competing constraints at once, then there are likely to be only a small number of ways to solve it. So we expect different systems that solve the same hard problem (for example, neural networks and the brain) to converge to the same solution (see also Huh et al.,, 2024; Hosseini et al., 2024a, ). If a problem is relatively simple, then we might expect many different solutions to work. For instance, there are famously scores of algorithms that have been proposed for sorting numbers in an array.
@@
Vision is not like sorting a list of numbers: it is a hard problem that requires satisfying many constraints (e.g., fast processing, reliable transmission of input, invariance to different light conditions, among many others). Therefore, we can expect that any system that solves the problem will share core features with any other. We see this not only in the comparison of humans vs. neural networks, but also in the comparison of humans vs. monkeys (Rajalingham et al.,, 2018), and primates vs. animals whose visual cortices arise from totally different evolutionary phylogenies, such as cephalopods (Pungor et al.,, 2023).
@@
The problem of learning and using language is also not like sorting a list of numbers: an entity that learns and processes language has to satisfy many constraints (storage of lexical items, generalization to novel contexts, fast processing, etc.). And as we have discussed, large Transformer-based models do learn and deploy nontrivial aspects of language. If the Contravariance Principle is right, then it is reasonable to expect some degree of commonality in the solutions that exist in the brain and in Transformers. Indeed, the massive literature on linguistic interpretability in Transformers, which we partially reviewed in Section 2.2, has revealed representational strategies in neural networks that lead to correct novel predictions about human performance (Lakretz et al.,, 2021), and the way that Transformers process syntactic features like agreement cues has close parallels with independently-proposed cognitive frameworks for modeling human language processing based on cue-based retrieval (Lewis and Vasishth,, 2005; Ryu and Lewis,, 2021; Timkey and Linzen,, 2023).
@@
Thus, it is just not the case that these models can be dismissed out of hand because they are different from the brain. To do so would be to miss out on a source of in silico insight and inspiration which has been hugely useful in other areas of cognition.
@@ 3.2.1 The significance of the learning problem in linguistics and cognitive science
For a century, cognitive scientists, linguists, computer scientists, and philosophers developed formal theories of how language learning—or any learning—must work. The key concept has been what knowledge is brought to the learning process by a learner beyond the data, known as inductive bias in the machine learning literature (Mitchell,, 1980; Goyal and Bengio,, 2022), as illustrated in Figure 3.2.1. We can think of a learner as a device that takes in some data and outputs a hypothesis, or a set of hypotheses, or a probability distribution over hypotheses, for the underlying process generating the data. For example, given a bunch of data points in a 2D plane, one hypothesis would be that the points are generated along a line in the plane; this is the assumption underlying linear regression. Given linguistic utterances as data, one might hypothesize a certain grammar underlying the data.
The key issue is that, in a very general sense, the data fundamentally underdetermine the hypothesis that the learner arrives at, because any data is compatible with many, many hypotheses. In Figure 3.2.1A, where the learning data is points and the hypotheses are lines, both the straight and the curvy lines fit the learning data equally well; the fact that the straight line seems like the right hypothesis—and is likely the one that humans would seize on—therefore cannot be a function of fit to the data. It must be a function of fit to the data plus something that biases a learner to favor one hypothesis over another, even among hypotheses that fit the data equally well: this is inductive bias. Inductive bias is something that the model (or modeler) brings to the problem, not something inherent in the data. [^6] We use the term inductive bias here to refer to anything that causes a learner to prefer one hypothesis over another, beyond any function of input data (Mitchell,, 1980). Such biases exist and are necessary even in learning algorithms that are not strictly ‘inductive’, and they may or may not be identifiable as Bayesian priors.
@@ B. Linguistic data
(a) Illustration of the role of inductive bias in generalization. A. The black points represent data, and the lines represent generalizations that could be formed on the basis of the data. Both lines fit the data exactly, but make different predictions for new datapoints. Intuitively, the blue line seems like a better hypothesis. But the choice to prefer any single generalization over any other can only be a function of inductive bias—a preference for hypotheses which are in some sense ‘simpler’—which is not a function of the data. B. Inductive bias in language. In black, possible corpus data from English. In blue, sentences which generalize from the corpus data on the basis of hierarchical structure. In red, sentences which generalize on the basis of linear word order. Intuitively, the blue sentences seem like natural generalizations, and the red ones seem unnatural, even though all are equally consistent with the data given. This phenomenon is claimed to be the result of innate biases specific to language (Chomsky,, 1971).
@@
Inductive bias shows up in several forms and under many names in the scientific literature. In linguistic theory, it appears perhaps first as the evaluation procedure in Chomsky, (1965, pp. 30–48), a function which compares two different grammars that can generate an observed set of sentences (that is, two grammars that fit some learning data equally well), and ranks them in order of preference, in a way that captures how humans generalize beyond the data (Chomsky,, 1965, p. 45). More generally, it appears as Universal Grammar (UG)—an “innate schematism of mind that is applied to the data of experience” (Chomsky,, 1971, p. 28) that enables language learning and generalization; furthermore, UG is held to be domain-specific to language (not applying in to any other aspect of cognition) and species-specific to humans (Huybregts,, 2019). In this approach, (generative) grammatical theories such as Minimalism are hypotheses about the nature of UG (Chomsky,, 1993; Adger,, 2003). They are meant to precisely delimit what languages may exist and be acquired by humans.
@@
This approach to linguistic theory serves the goal of explanatory adequacy: the idea is that a theory of grammar should not only capture which sentences are grammatical and ungrammatical in a particular language, but also that the theory encompass all and only the possible languages that we might actually find in the world (Chomsky,, 1965, Ch. 1). Thus, if a grammatical theory seems to accommodate ‘impossible’ or unattested kinds of languages, then this would be considered evidence against that theory, because the theory is supposed to represent the universal and language-specific “schematism of mind” that a learner applies to language data. [^7] For example, under this view, it would be justified to criticize a theory of grammar such as Head-Driven Phrase Structure Grammar (HPSG: Pollard and Sag,, 1994; Sag et al.,, 2003) on the basis that it is Turing-complete, capable of generating any recursively enumerable formal language. This approach to explaining linguistic phenomena has been called “explanation by constrained description” (Haspelmath,, 2009, pp. 384–385).
@@
Naturally then, Universal Grammar has been the major focus of study in language learning, seen as the simultaneous solution to two different problems: (1) how children can learn language from inadequate data (the classic Argument from the Poverty of the Stimulus: see Pearl,, 2022, for a review), and (2) why human language is the way it is: because Universal Grammar strongly restricts the set of possible languages. [^8] Given the centrality of this reasoning to much of linguistic theory, one might expect it to be backed up with strong experimental evidence that humans cannot or do not learn languages which violate the putatively universal principles of human languages. This is not the case: there is only limited and ambiguous experimental evidence for hard formal limits on human language learning. Among the most on-target existing studies is Smith et al., (1993), whose object of study was a man described as a polyglot savant living in a mental health facility. This individual and four control subjects (linguistics undergraduates) were tasked with learning artificial languages designed to be ‘impossible’ in three ways: (1) negation and tense are indicated by word order, (2) there is an agreement pattern judged to be impossible, and (3) the position of an emphatic marker is determined by a rule involving counting words. Results are not systematically reported, but seem to indicate that the polyglot was able to learn the ‘impossible’ word order and agreement rules (1) and (2), but not the rule for the emphatic marker (3). Another commonly cited experimental work on this topic is Musso et al., (2003), who expose German speakers to Italian and Japanese sentences, either following the real rules of those languages, or following modified rules deemed to be linguistically impossible, for example placing a negation marker after the third morpheme from the beginning of a sentence. The result is equally accurate learning of the ‘natural’ and ‘unnatural’ languages. fMRI on the subjects shows that the real languages elicit activity in the left inferior frontal gyrus, while the unnatural ones elicit activity elsewhere. We believe the meaning of these results is unclear. Only a small number of languages and participants (all of whom were already native speakers of largely hierarchically-structured languages) were tested, the localization of syntax in the brain is still contentious, and the patterns of brain activity for the ‘unnatural’ languages might reflect a lack of practice with such patterns, rather than their impossibility or a qualitative difference between linear and hierarchical rules. The dream has been to come up with a formalism for linguistic description which captures human generalizations, is domain-specific to language, and arises from a genetic endowment unique to humans (Chomsky,, 1988; Hauser et al.,, 2002; Berwick et al.,, 2011)—thus in one fell swoop solving (1) and (2).
@@
Under this logic, it follows that a good theory of language learning should be restrictive: that is, there should be languages that cannot be learned under the theory, and this restriction on the hypothesis space provides explanatory adequacy. [^9] See, for example, Kodner et al.,’s (2022) criticism of Yang and Piantadosi,’s (2022) model of language learning as Bayesian program induction, a model which successfully learns grammars of various formal classes given small amounts of string input, thus addressing the Poverty of the Stimulus problem for formal structures. Although the model readily meets the challenge of inducing formal structure from strings, it has been dismissed by some in the linguistics literature because the same model could also learn grammars that are unlike human language. In terms of Figure 3.2.1, the intuitively bad hypotheses would be ruled out as simply unavailable as mental representations during learning (Everaert et al.,, 2015). [^10] Although these are common examples, it is not clear exactly how generative formalisms rule out the unnatural hypotheses here. In particular, the languages implied by the unnatural hypotheses are context-free, just as much as the languages implied by the natural hypotheses. So these (string) languages could be generated from, for example, Minimalist Grammars (Chomsky,, 1993; Stabler,, 1997), since Minimalist Grammars generate a superset of context-free languages (Michaelis,, 1998). This approach to linguistic explanation has a pleasing elegance to it: learners must be restricted to learn properly, and we see that the variation in actual languages is restricted, therefore we can kill two birds with one stone by finding the right representational restrictions on learners.